Janet rides her bike to work at a speed of 12 miles per hour. It takes her 20 minutes to get there from her house. How long is her ride to work?
Answer by ReadingBoosters(3246)

20 min x 12 m/h x 1/60 h/min leaves just miles as units
20 x 12 x 1/60 = 240/60 = 4 miles